OpenStudy (anonymous):

cot(3x)= squareroot of 3 what is the solution set

OpenStudy (cwrw238):

what do you get ? you'll get an error or similar message since angles can only of cosine detween -1 and 1

OpenStudy (campbell_st):

you should look at using cot = 1/tan so take the reciprocals of both sides and you get \[\tan(3x) = \frac{1}{\sqrt{3}}\] then find the arc tan \[3x = \tan^{-1}(\frac{1}{\sqrt{3}})\] hint... its an exact value angle from the 30, 60 right triangle when you get the answer, divide it by 3 to find x hope this helps

OpenStudy (cwrw238):

oops sorry - i misread the question!